OpenStudy (anonymous):

Which of the following expressions are polynomials select all correct answers A. 1+1/X b.0 c.-XY-Z d.y^x

OpenStudy (anonymous):

c. and d. a polynomial is a variable function which has positive values. n hey, dont forget the medal (y)

OpenStudy (anonymous):

A polynomial has the general form \[P(x)=a_nx^n+a_{n-1}x^{n-1}+\cdots+a_2x^2+a_1x+a_0\] where \(a_i\in\mathbb{C}\) for \(i=0,1,2,...,n\). Notice that \(n\) is a non-negative integer. (a) can't be a polynomial because \[1+\frac{1}{x}=1+x^{-1}\] which has a negative exponent on one of the terms. (b) is a polynomial - it consists of only a constant term, \(a_0\). (c) can't be a polynomial because there is more than one variable. However, if \(y\) and \(z\) are kept fixed, you have linear expression: \[-xy-z=a_1x^1+a_0\] so it could be considered a polynomial. Similarly, if \(y\) is the variable and \(x\) and \(z\) are constant, you have another linear polynomial; the same is true if \(x\) and \(y\) are constant and \(z\) is the variable. (d) is not a polynomial because there are multiple variables. However, if you fix \(x\) as a non-negative integer, you have a polynomial in terms of \(y\), like \(y^2\) or \(y^{10}\). Fixing \(y\) doesn't yield a polynomial, though.